Find the center of mass (gravity) of a solid of constant density that is bounded by the parabolic
cylinder x = y
2
and the planes x = z, z = 0, and x = 1. (
1
Expert's answer
2020-04-03T15:53:59-0400

The solid V and its projection on the "xy-"plane are given in Figures 01 and 02.



The lower and upper surfaces of the "V" are planes "z=0" and "z=x." Describe "V" as a region


"V=[(x,y,z):-1\leq y\leq 1, y^2\leq x\leq 1,0\leq z\leq x]"

If the density function "f(x, y, z)=k, k=const," then the mass is


"m=\iiint_VfdV=\displaystyle\int_{-1}^1\displaystyle\int_{y^2}^1\displaystyle\int_{0}^xkdzdxdy="

"=k\displaystyle\int_{-1}^1\displaystyle\int_{y^2}^1(x-0)dxdy=k\displaystyle\int_{-1}^1\bigg[{x^2\over 2}\bigg]\begin{matrix}
 1 \\
 y^2
\end{matrix}\ dy="


"={k\over 2}\displaystyle\int_{-1}^1(1-y^4)dy={k\over 2}\bigg[y-{y^5\over5}\bigg]\begin{matrix}
 1 \\
 -1
\end{matrix}="

"={k\over 2}(1-{1\over 5}+1-{1\over 5})={4k\over 5}"

Consider a function which is bounded by the curves "z=g_1(x, y)"  and "z=g_2(x, y)"


"g_1(x, -y)=g_1(x, y), \ g_2(x, -y)=g_2(x, y)"

Then we say that the region is symmetric about "y-" axis and hence "M_{xz}=0," and the "y-"coordinate of the center of mass is zero "\bar{y}=0."


"M_{yz}=\iiint_VxfdV=\displaystyle\int_{-1}^1\displaystyle\int_{y^2}^1\displaystyle\int_{0}^xkxdzdxdy="

"=k\displaystyle\int_{-1}^1\displaystyle\int_{y^2}^1x(x-0)dxdy=k\displaystyle\int_{-1}^1\bigg[{x^3\over 3}\bigg]\begin{matrix}
 1 \\
 y^2
\end{matrix}\ dy="

"={k\over 3}\displaystyle\int_{-1}^1(1-y^6)dy={k\over 3}\bigg[y-{y^7\over7}\bigg]\begin{matrix}
 1 \\
 -1
\end{matrix}="

"={k\over 3}(1-{1\over 7}+1-{1\over 7})={4k\over 7}"

"M_{xy}=\iiint_VzfdV=\displaystyle\int_{-1}^1\displaystyle\int_{y^2}^1\displaystyle\int_{0}^xkzdzdxdy="

"={k\over 2}\displaystyle\int_{-1}^1\displaystyle\int_{y^2}^1x^2 dxdy={k\over 2}\displaystyle\int_{-1}^1\bigg[{x^3\over 3}\bigg]\begin{matrix}
 1 \\
 y^2
\end{matrix}\ dy="

"={k\over 6}\displaystyle\int_{-1}^1(1-y^6)dy={k\over 6}\bigg[y-{y^7\over7}\bigg]\begin{matrix}
 1 \\
 -1
\end{matrix}="

"={k\over 6}(1-{1\over 7}+1-{1\over 7})={2k\over 7}"

Therefore, the centre of mass is


"(\bar{x}, \bar{y},\bar{z})=\big({M_{yz}\over m},{M_{xz}\over m},{M_{xy}\over m}\big)=\big({5\over 7},\ 0,\ {5\over 14}\big)"